Good evening, I have the problems about the GP and Ap. Please help me in.....

1) A ball is dropped from a height of 18 ft. If on it rebound, it rises to a height 2/3 the distance from which it fell, how far ( up and and down ) will the ball have traveled when it hits the ground for the sixth times? (Give the approximately answer)
2) A pendulum is released and swings through an arc measuring 15 inche. Each swing thererater it travels a distance 0.85 times the length of the previous pass. Approximately how far will the pendulum swing on the eighth pass?
3) A readioactive dye is injected into a system in a medical test. After one hour, 60% of the dye remains. At the end of two hours 60% of the remaining dye remains, and so on. If one unit of dye is injected, approximately what percent will remain after 12 hours?

Answer by venugopalramana(3286)

1) A ball is dropped from a height of 18 ft. If on it rebound, it rises to a height 2/3 the distance from which it fell, how far ( up and and down ) will the ball have traveled when it hits the ground for the sixth times? (Give the approximately answer)
HT. DURING I DROP = H1 = 18'
HT DURING II DROP = H2 = (2/3)*H1=2*18/3=12
HT.DURING III DROP = H3 = (2/3)*H2=[(2/3)^2]H1
.........................................
..........................................
HT DURING N TH. DROP = HN = [(2/3)^(N-1)]H1
IT HIT GROUND 6 TIMES..HENCE WE WANT HT FOR THE 7 TH. DROP.
THAT IS FOR N=7 WE HAVE H7 = 18=1.58'


2) A pendulum is released and swings through an arc measuring 15 inche. Each swing thererater it travels a distance 0.85 times the length of the previous pass. Approximately how far will the pendulum swing on the eighth pass?
PROCEEDING IN THE SAME MANNER AS ABOVE,WE GET
SN=15